# MatFormer: Nested Transformer for Elastic Inference

### v1

@@ Appendix C Scaling Laws for Language Decoders
We provide results split by granularities for validation loss, average score on RANK tasks, average score on GEN tasks, and consistency in Figures 7, 8, 9, and 10 respectively. We observe that while the gap in validation loss between MatLMs and Baselines appears to be constant, the gap for downstream evaluations reduces with scale - in fact, granularities L, M and S have better downstream performance for models larger than 1B. For consistency, the gap appears to reduce with scale, but one would need to scale the models by many orders of magnitude beyond what’s possible today for baselines to have comparable consistency with MatLMs.
C.1 Scaling laws of MatFormers vs Transformers.
@@
Scaling laws are essential tools to estimate optimality under as the cost of training or inference is increased. Scaling laws can take diverse viewpoints such as overall training cost in FLOPS, training data and parameter efficiency, and inference mean FLOPS utilization vs latency for deployments.
The scaling relationship of MatFormers versus Transformers is both simple and complex. Simple, because MatFormers scaling curves for pretraining are only slightly offset from Transformers – thus MatFormers only require a fixed relative amount of additional compute and the same hyperparameters that work for Transformers are effective for MatFormers. For the setting where we use the same hyperparameters as Transformers, MatFormers need at most $10-20\%$ more training tokens to reach the same loss as a regular Transformer. Initial experiments where we tune hyperparameters for the individual forward/backward passes and by performing more careful initialization of the subslices the gap appears to shrink. While we do not have enough data to make definite statements, it appears MatFormer scaling can be improved to be close to Transformers scaling needing less than $0-5\%$ additional training tokens.
@@
Appendix D Further Analysis on Language Decoders
D.1 KL Divergence Between S, M, L and XL Models
@@
Figure 6: The smoother variant of consistency measures the KL divergence between the smaller models and the corresponding XL model. This metric, unlike the exact match accuracy variant, also accounts for different sampling strategies on the output distribution during deployment. In this figure, we plot KL divergence of S, M, L granularities with respect to XL for the 2.6B parameter model.
D.2 Ablations on Training Method
We experiment with several aspects of our training method on a 850M parameter MatLM. Our training procedure is unique compared to others (further discussed in Section 2) in 2 ways: (a) we learn all granularities in the same weight space and (b) we use joint optimization as described in Section 3. To assess the effect of these differences on performance, first we train a Transformer model with independent FFN modules with {S, M, L, XL} granularites using joint optimization (Independent modules). Next, we train a MatLM model with the only difference being that at each step, we optimize for a single granularity chosen uniformly at random (Subsampling). We find that joint optimizing a MatLM performs significantly better than these baselines, implying efficacy of both aspects of our training method.
Table 5: We compared the validation loss of models from Joint Optimization to training MatLMs with independent MLP modules for each granularity (Independent modules) and sampling a single granularity to optimize for at each step (Subsampling) for 850M parameter models. We find that Joint Optimization performs significantly better than both these methods.
Model Training Strategy XL L M S
Baseline - 2.840 2.910 2.9710 3.017
MatFormer Joint Optimization 2.874 2.928 2.980 3.030
Independent MLP modules 2.894 2.942 2.985 3.030
Subsampling 2.929 2.946 2.999 3.049
We discuss additional ablations such as re-weighting losses to improve the performance of the XL model in Appendix D.4, and additionally studied scaling trends for these ablations. We found the reweighting loss trick to be especially powerful, bringing the performance on downstream evals within $0.1\%$ for the XL model. This also nudges us towards finding better hyperparameters and weight initializations for reliable scaling of MatLMs (Yang et al., 2022).
D.3 Changing Embedding Size
Because of the ubiquity of 64k vocabs size (Brown et al., 2020) we additionally train models upto 201M non-embedding parameters similar to those described in Appendix A, except that the embedding size is 64k (the largest model corresponds to the 463M parameter model). We plot the scaling trends in Figure 11. Though 4 models is not enough to extrapolate a trend, we observe that the scaling trend for validation loss appears to be similar.
D.4 Reweighting Strategies
We additionally experiment with reweighting the losses for the individual granularities in order to boost the performance of the largest granularity while minimally impacting the performance of the smaller granularities. We present the relative weights used in Table 6 as $\lambda_{4}:\lambda_{3}:\lambda_{2}:\lambda_{1}$, and find that in general, upweighting the largest granularity greatly improves quality. Another interesting related direction for improving MatFormer performance further is granularity appropriate initialization (Yang et al., 2022).
@@ D.4 Reweighting Strategies
Table 6: For 850M model, we experiment with modifying $\mathcal{L}_{\textsc{joint}}$ to use a weighted average as opposed to an unweighted average, and report the results across all granularities. We find that all strategies that upweight the loss for the largest granularity perform well, with modest degradation on the M and S granularties.
Model Relative Weights XL L M S
Baseline N/A 2.840 2.910 2.971 3.017
MatFormer 1:1:1:1 2.874 2.928 2.980 3.030
$2:1.5:1.25:1$ 2.867 2.927 2.986 3.051
$1:1.25:1.5:2$ 2.883 2.936 2.982 3.026
$2:1:1:1$ 2.863 2.929 2.985 3.043
$\sqrt{8}:\sqrt{4}:\sqrt{2}:1$ 2.862 2.924 2.990 3.063
D.5 Scaling Laws for Reweighted Strategy
We conduct scaling experiments similar to those described in Section 4.1 for the reweighed models, specifically for models with the ratio $2:1.5:1.25:1$, and plot the results in Figure 12. We note that the scaling trend is similar to the MatLM with a $1:1:1:1$ relative weighting ($a=19.889,b=-0.130,c=1.374$), but with a slightly better validation loss .
Appendix E Further Analysis on Vision Encoders
E.1 Decoupling Effect of MatFormer on Pretraining and Finetuning
@@
Table 7 investigates the effect of MatFormer on pretaining and finetuning phases of ViT-L/16 model. ViT-L/16 is typically pretrained on ImageNet-21K and then finetuned on ImageNet-1K for the final evaluation. Table 7 shows that having a MatFormer during pretraining generates a better model for downstream finetuning compared to regular ViT pertaining. At the same time, finetuning a vanilla pretrained ViT with MatFormer results in flexibility being induced into the model. Despite being up to $2\%$ less accurate than its counterparts at some granularities, a fine-tuned MatViT learned to reallocate the information to provide strong nested models. Considering that this is insignificant compared to pretaining costs, possible to take the largest pretrained ViT model and finetune with MatFormer to obtain a deployable MatViT variant.
@@
Table 7: 2 × 2 grid of pairs to evaluate the effects of MatFormer and standard training on the pretraining (PT) on ImageNet-21K and finetuning (FT) on ImageNet-1K using a L/16 architecture. Using a MatFormer during pretraining helps bring more accurate, and elastic encoders for downstream uses.
@@
E.2 Traditional Image Retrieval Evaluation
Table 8 showcases traditional image retrieval evaluation on ImageNet-1K where the query and the document encoders are the same for nearest neighbor retrieval. The 1-nearest neighbor (NN) based evaluation closely follows one-vs-all classification results shown in Figure 4. Both MatViT variants B/16 and L/16 have submodels that have as good or better retrieval performance compared to their independently trained counterparts. Concretely, MatViT-based retrieval can be up to $0.5\%$ more accurate than the baselines while a 200M parameter MatViT submodel can be more accurate than the 300M parameter ViT baseline.
@@ E.2 Traditional Image Retrieval Evaluation
Figure 7: We train various decoder-only MatLM models at a range of sizes from 78M to 2.6B parameters and observe the scaling trends for each model granularity on validation loss. We observe that the gap between MatLM and the baseline appears to be constant at each granularity. The consistency between the submodels of granularities and the XL models shows the effect of MatFormer joint training on natively ensuring similar behavior across submodels.
(a) XL-model Average Score on GEN tasks
(b) L-model Average Score on GEN tasks
(c) M-model Average Score on GEN tasks
(d) S-model Average Score on GEN tasks
(e) Average Score on GEN tasks for all granularities - XL, L, M, S.
Figure 8: We train various decoder-only MatLM models at a range of sizes from 78M to 2.6B parameters and observe the scaling trends for each model granularity for the average score on GEN tasks 1-shot evaluation. We observe that the gap between MatLM and the baseline reduces with scale, outperforming the baselines for S, M, L granularities for the largest models.
(a) XL-model Average Score on RANK Evals
(b) L-model Average Score on RANK Evals
(c) M-model Average Score on RANK Evals
(d) S-model Average Score on RANK Evals
(e) Average Score on RANK Evals for all granularities - XL, L, M, S
Figure 9: We train various decoder-only MatLM models at a range of sizes from 78M to 2.6B parameters and observe the scaling trends for each model granularity for the average score on RANK 1-shot evaluation. We observe that the gap between MatLM and the baseline reduces with scale, outperforming the baselines for S, M, L granularities for the largest models.
@@
Figure 10: We train various decoder-only MatLM models at a range of sizes from 78M to 2.6B parameters and observe the scaling trends for each submodel S, M, L for the consistency with the XL model. We observe that the gap between MatLM and the baseline reduces with scale, but one would need to scale the baseline by many orders of magnitude to have consistency comparable to that of MatLMs.
(a) XL-model Loss
(b) L-model Loss
(c) M-model Loss
(d) S-model Loss
(e) Loss for all granularities - XL, L, M, S
Figure 11: We train various decoder-only MatLM models at a range of sizes from 29M to 267M parameters with an embedding size of 64k and observe the scaling trends for each model granularity on validation loss. We observe that the gap between MatLM and the baseline appears to be constant at each granularity, similar to what is observed in Figure 7.
(a) S-model Loss
(b) S-model Loss
(c) S-model Loss
(d) Loss for all granularities - XL, L, M, S
(e) All
Figure 12: We train various decoder-only MatLM models at a range of sizes from 78M to 2.6B parameters with a reweighing ratio of $2:1.5:1.25:1$ and observe the scaling trends for each model granularity on validation loss. We observe that the gap between MatLM and the baseline appears to be constant at each granularity, similar to what is observed in Figure 7.
@@
TriviaQA (EM) 0.14 0.16 0.19 0.25 0.14 0.3 0.19 0.28
NaturalQuestions (EM) 0.06 0.03 0.03 0.06 0.03 0.03 0.03 0.03
WebQuestions (EM) 0.1 0.2 0.15 0.2 0.2 0.3 0.3 0.3
LAMBADA 0.06 0.02 0.02 0 0.02 0 0 0
HellaSwag 25.42 26.28 26 25.87 25.95 25.9 25.95 25.94
StoryCloze 52.81 53.39 53.13 53.34 54.46 53.5 54.46 54.36
WSC 52.98 51.93 53.68 50.88 55.79 54.04 52.28 52.63
WinoGrande 48.46 51.54 51.54 47.99 50.99 48.46 48.86 49.41
Winograd 53.11 52.75 52.38 53.85 55.31 55.31 52.75 55.68
SQuAD v2 (EM) 11.19 36.71 33.14 33.77 20.08 29.17 22.78 30.97
RACE-H 25.53 25.84 24.73 25.44 26.07 25.9 25.96 25.84
RACE-M 29.18 30.15 28.83 29.94 28.83 30.43 29.74 31.48
PIQA 55.77 55.22 54.62 55.28 54.52 54.79 56.86 54.08
ARC-C 21.5 20.9 21.08 21.67 21.59 21.33 22.35 22.1
ARC-E 34.55 35.48 34.3 35.73 34.89 36.11 34.55 35.98
OpenBookQA 25.4 28.6 27.6 28 28.2 28 29.8 29
BoolQ 48.72 44.89 51.87 47.37 51.28 46.85 52.11 45.87
COPA 62 64 62 61 63 63 60 60
RTE 53.79 52.35 52.35 51.99 51.26 54.51 51.99 52.71
WiC 49.53 47.34 49.06 47.34 47.34 47.34 47.65 47.34
MultiRC (F1) 53.17 51.72 53.42 53.28 56.86 53.82 55.46 53.42
ReCoRD 39.52 39.22 40.03 39.95 40.55 40.42 40.8 40.83
CB 41.07 42.86 44.64 39.29 44.64 41.07 42.86 44.64
ANLI-R1 30.9 32 32.3 31.9 32.5 32.3 32.5 31.7
ANLI-R2 31.1 30.9 31.1 30.1 30.7 30.8 30.6 30.3
ANLI-R3 31.75 30.75 30.58 30.25 30.33 29.67 30 30.17
Average 33.76 34.82 34.95 34.41 34.83 34.74 34.65 34.81
Avg over GEN Taks 2.31 7.42 6.7 6.85 4.09 5.96 4.66 6.31
Avg over RANK Tasks 41.25 41.34 41.68 40.97 42.15 41.6 41.79 41.59
Dev set log pplx 4.010 4.012 3.97 3.96 3.905 3.908 3.83 3.868
@@
TriviaQA (EM) 1.04 0.9 0.98 1.26 1.16 1.89 1.86 2.00
NaturalQuestions (EM) 0.08 0.11 0.14 0.08 0.3 0.11 0.28 0.11
WebQuestions (EM) 0.59 0.94 0.44 0.98 1.28 0.89 1.33 0.79
LAMBADA 0.16 0.68 0.43 1.16 1.51 0.95 0.49 0.99
HellaSwag 27.77 27.3 27.45 27.61 27.58 27.84 28.86 28.56
StoryCloze 56.33 56.07 57.03 56.87 57.3 57.78 58.63 58.52
WSC 55.44 55.44 56.49 60.35 58.25 58.6 57.54 58.6
WinoGrande 52.01 50.12 50.28 49.17 51.22 50.43 51.54 49.09
Winograd 54.21 55.68 56.78 57.51 61.54 58.61 60.44 61.17
SQuAD v2 (EM) 22.13 17.28 20.05 18.02 26.42 11.42 25.76 16.53
RACE-H 27.93 27.9 27.5 28.53 28.7 28.82 28.73 28.73
RACE-M 33.29 34.47 34.19 34.05 34.54 33.91 33.29 34.19
PIQA 57.13 58.05 56.91 57.94 57.94 58.00 59.52 58.92
ARC-C 22.53 22.61 23.63 22.27 24.06 22.1 24.66 23.55
ARC-E 40.24 39.39 40.19 40.49 41.71 40.74 41.62 41.16
OpenBookQA 30.60 31.00 30.80 31.80 31.00 32.80 34.00 32.6
BoolQ 54.13 52.23 52.45 52.05 55.63 52.17 55.9 48.44
COPA 62 61 61 61 61 64 64 65
RTE 52.71 53.07 52.35 53.43 50.54 52.71 52.71 52.71
WiC 47.34 51.41 47.34 49.37 47.96 47.81 47.65 47.34
MultiRC (F1) 54.34 53.34 45.65 56.12 47.47 52.62 47.62
ReCoRD 48.58 49.4 48.99 50.13 50.56 51.25 52.82 52.51
CB 42.86 44.64 42.86 44.64 39.29 44.64 42.86 42.86
ANLI-R1 31.8 32.6 31.8 32.4 32.4 32.8 32.2 32.1
ANLI-R2 30.5 29.8 31.1 29.8 32.00 30.5 30.5 30.1
ANLI-R3 30.08 30.25 30.5 32.00 33.5 31.42 30.67 30.42
Average 35.99 35.51 35.96 36.1 37.06 36.14 37.33 36.33
GPT3-GEN 4.8 3.98 4.41 4.3 6.14 3.05 5.94 4.08
GPT3-RANK 43.42 43.02 43.48 43.67 44.42 44.02 44.8 44.01
Dev set log pplx 3.55 3.55 3.512 3.505 3.456 3.458 3.354 3.40
@@
TriviaQA (EM) 2.09 2.4 2.2 3.17 2.84 2.73 5.18 3.12
NaturalQuestions (EM) 0.11 0.28 0.28 0.5 0.58 0.3 0.91 0.61
WebQuestions (EM) 2.12 1.38 1.08 1.67 1.67 1.43 2.41 1.57
LAMBADA 0.29 1.79 0.66 1.92 1.9 2.46 2.76 2.64
HellaSwag 29.89 29.69 30.05 30.02 31.18 30.63 32.52 31.58
StoryCloze 59.17 58.85 59.54 60.13 60.24 60.5 61.68 61.36
WSC 61.05 59.65 59.3 58.6 61.75 56.84 58.95 57.19
WinoGrande 51.46 52.88 49.57 50.91 52.41 50.75 50.91 52.01
Winograd 55.68 56.04 57.88 59.71 63 59.71 61.17 60.07
SQuAD v2 (EM) 22.38 22.79 13.38 17.83 20.03 18.66 22.03 21.81
RACE-H 29.45 28.33 28.9 28.67 29.22 29.07 29.67 28.79
RACE-M 35.31 36.14 36.14 36.91 36.42 36.14 37.6 36.07
PIQA 58.98 59.9 59.58 59.85 59.79 60.45 62.19 60.61
ARC-C 23.38 20.82 23.21 21.33 23.81 23.21 25 22.95
ARC-E 42.3 42.34 44.11 43.52 44.53 44.44 46.8 45.62
OpenBookQA 32.8 35.2 34.6 36.4 35.2 35.8 36.8 36.6
BoolQ 53.43 59.05 55.32 58.72 52.87 57.22 54.22 55.6
COPA 61 61 61 66 64 63 60 66
RTE 52.71 54.51 53.43 51.62 51.62 53.07 54.15 49.46
WiC 47.18 48.43 47.65 49.22 47.65 50.16 47.34 51.25
MultiRC (F1) 53.07 51.69 53.5 51.36 48.46 47.14 45.72 46.23
ReCoRD 54.34 53.86 55.18 55.33 56.75 56.79 58.39 58.07
CB 42.86 46.43 42.86 46.43 42.86 46.43 50 51.79
ANLI-R1 32 31.3 32 32.2 32.5 32.3 32.2 32.8
ANLI-R2 32.6 30.2 30.9 29.8 30.6 31.2 29.8 30.9
ANLI-R3 32.08 29.25 30.75 30.08 32.17 31.25 31.5 32.17
Average 37.22 37.47 37.04 37.77 37.85 37.76 38.46 38.34
Avg over GEN Taks 5.4 5.73 3.52 5.02 5.41 5.12 6.66 5.95
Avg over RANK Tasks 44.8 45.03 45.02 45.56 45.57 45.53 46.03 46.05
Dev set log pplx 3.31 3.33 3.30 3.285 3.224 3.235 3.15 3.18

### v2

@@ Appendix E Scaling Laws for Language Decoders
We provide results split by granularities for validation loss, average score on evaluation tasks, and consistency in Figures 9, 10, and 11 respectively. We observe that the gap in validation loss and evaluation accuracy between MatLMs and Baselines appears to be constant. For consistency, the gap appears to reduce with scale, but one would need to scale the models by many orders of magnitude beyond what’s possible today for baselines to have comparable consistency with MatLMs.
E.1 Scaling laws of MatFormers vs Transformers.
@@
Scaling laws are essential tools to estimate quality as the cost of training or inference is increased. Scaling laws can help us consider various nuances of training and deployment such as overall training cost in FLOPS, training data and parameter efficiency, and inference mean FLOPS utilization vs latency for deployments.
@@
Appendix F Further Analysis on Language Decoders
F.1 KL Divergence Between S, M, L and XL Models
@@
Figure 7: The smoother variant of consistency measures the KL divergence between the smaller models and the corresponding XL model. This metric, unlike the exact match accuracy variant, also accounts for different sampling strategies on the output distribution during deployment. In this figure, we plot KL divergence of S, M, L granularities with respect to XL for the 850M parameter model.
F.2 Using MatFormers for Attention Sub-Block
We experiment with applying MRL to the attention sub-block of the Transformer block. More specifically, we jointly optimize 4 subnetworks with a varying number of attention heads $n_{attn}$, $43*n_{attn}/$ 4, $n_{attn}/2$, $n_{attn}/4$ of size $d/n_{attn}$ each. In Figure 8, we plot the validation loss for these models (MatLM-Attn), their corresponding baselines and Mix’n’Matched subnetworks. We find that similar to our experiments on Mix’n’Matched MatLMs, Mix’n’Match helps obtain many MatLM-Attn models on the optimal loss-vs-compute curve.
(a) Validation loss
Figure 8: Validation loss for the 850M MatLM-Attn & baseline models.
@@ F.2 Using MatFormers for Attention Sub-Block
Table 6: For 850M model, we experiment with modifying $\{p_{S},p_{M},p_{L},p_{XL}\}$ to sample submodels from a non-uniform distribution during training and report the results across all granularities. We find that all strategies that upweight the loss for the largest granularity perform well, with modest degradation on the M and S granularties.
Model Probabilities S M L XL
Baseline N/A 3.017 2.971 2.910 2.840
MatFormer 0.44/0.31/0.15/0.10 2.963 2.925 2.899 2.877
0.31/0.27/0.22/0.20 2.977 2.929 2.890 2.857
0.25/0.25/0.25/0.25 2.970 2.934 2.886 2.846
0.20/0.22/0.24/0.34 3.000 2.939 2.885 2.836
0.17/0.20/0.22/0.41 3.010 2.943 2.884 2.829
F.3 Tuning Sampling Probability
In Table LABEL:table:reweigh, we experiment with tuning the sampling probabilities for individual granularities in order to investigate the trade-off between granularity size and performance. More specifically, we tune $\{p_{1},p_{2},p_{3},p_{4}\}$, and find that all strategies that upweight the loss for the largest granularity perform well, with modest degradation on the M and S granularties.
Appendix G Further Analysis on Vision Encoders
G.1 Decoupling Effect of MatFormer on Pretraining and Finetuning
@@
Table LABEL:tab:vit-2x2 investigates the effect of MatFormer on pretaining and finetuning phases of ViT-L/16 model. ViT-L/16 is typically pretrained on ImageNet-21K and then finetuned on ImageNet-1K for the final evaluation. Table LABEL:tab:vit-2x2 shows that having a MatFormer during pretraining generates a better model for downstream finetuning compared to regular ViT pertaining. At the same time, finetuning a vanilla pretrained ViT with MatFormer results in flexibility being induced into the model. Despite being up to $2\%$ less accurate than its counterparts at some granularities, a fine-tuned MatViT learned to reallocate the information to provide strong nested models. Considering that this is insignificant compared to pretaining costs, possible to take the largest pretrained ViT model and finetune with MatFormer to obtain a deployable MatViT variant.
@@
Table 7: 2 × 2 grid of pairs to evaluate (top-1 accuracy (%)) the effects of MatFormer and standard training on the pretraining (PT) on ImageNet-21K and finetuning (FT) on ImageNet-1K using a L/16 architecture. Using a MatFormer during pretraining helps bring more accurate, and elastic encoders for downstream uses.
@@
G.2 Traditional Image Retrieval Evaluation
Table LABEL:tab:ViT-KNN showcases traditional image retrieval evaluation on ImageNet-1K where the query and the document encoders are the same for nearest neighbor retrieval. The 1-nearest neighbor (NN) based evaluation closely follows one-vs-all classification results shown in Figure 4. Both MatViT variants B/16 and L/16 have submodels that have as good or better retrieval performance compared to their independently trained counterparts. Concretely, MatViT-based retrieval can be up to $0.5\%$ more accurate than the baselines while a 200M parameter MatViT submodel can be more accurate than the 300M parameter ViT baseline.
@@ G.2 Traditional Image Retrieval Evaluation
Figure 9: We train various decoder-only MatLM models at a range of sizes from 78M to 850M parameters and observe the scaling trends for each model granularity on validation loss. We observe that the gap between MatLM and the baseline appears to be constant at each granularity. The consistency between the submodels of granularities and the XL models shows the effect of MatFormer joint training on natively ensuring similar behavior across submodels.
(a) XL-model Average Score on Evals
(b) L-model Average Score on Evals
(c) M-model Average Score on Evals
(d) S-model Average Score on Evals
(e) Average Score on Evals for all granularities - XL, L, M, S
Figure 10: We train various decoder-only MatLM models at a range of sizes from 78M to 850M parameters and observe the scaling trends for each model granularity for the average score on 1-shot evaluation. We observe that the gap between MatLM and the baseline stays the same with scale, outperforming the baselines for on all granularities for the largest models.
@@
Figure 11: We train various decoder-only MatLM models at a range of sizes from 78M to 850M parameters and observe the scaling trends for each submodel S, M, L for the consistency with the XL model. We observe that the gap between MatLM and the baseline reduces with scale, but one would need to scale the baseline by many orders of magnitude to have consistency comparable to that of MatLMs.
@@
TriviaQA (EM) 0.14 0.13 0.19 0.14 0.14 0.11 0.19 0.10
NaturalQuestions (EM) 0.06 0.03 0.03 0.03 0.03 0.03 0.03 0.00
WebQuestions (EM) 0.10 0.10 0.15 0.10 0.20 0.15 0.30 0.15
LAMBADA 0.06 0.00 0.02 0.02 0.02 0.02 0.00 0.04
HellaSwag 25.42 26.59 26.00 26.28 25.95 25.95 25.95 26.32
StoryCloze 52.81 53.82 53.13 54.14 54.46 55.32 54.46 55.16
WSC 52.98 52.98 53.68 54.74 55.79 55.09 52.28 57.19
WinoGrande 48.46 47.91 51.54 50.36 50.99 51.38 48.86 50.20
Winograd 53.11 55.31 52.38 52.01 55.31 55.31 52.75 55.68
RACE-H 25.53 27.22 24.73 27.13 26.07 26.70 25.96 26.73
RACE-M 29.18 32.80 28.83 33.15 28.83 32.94 29.74 32.38
PIQA 55.77 56.47 54.62 56.31 54.52 56.42 56.86 56.09
ARC-C 21.50 22.01 21.08 21.50 21.59 21.76 22.35 21.76
ARC-E 34.55 34.55 34.30 34.68 34.89 34.97 34.55 34.55
OpenBookQA 25.40 28.00 27.60 29.00 28.20 30.60 29.80 30.60
BoolQ 48.72 52.14 51.87 52.08 51.28 52.14 52.11 52.20
COPA 62.00 58.00 62.00 59.00 63.00 58.00 60.00 59.00
RTE 53.79 53.43 52.35 53.07 51.26 52.35 51.99 52.35
WiC 49.53 47.49 49.06 47.18 47.34 47.34 47.65 47.18
MultiRC 51.28 53.38 51.51 53.96 47.67 53.57 49.38 52.74
RECORD 39.52 41.60 40.03 41.75 40.55 42.77 40.80 43.72
CB 41.07 41.07 44.64 41.07 44.64 41.07 42.86 42.86
ANLI-RI 30.90 32.60 32.30 32.30 32.50 32.40 32.50 32.70
ANLI-R2 31.10 30.60 31.10 30.60 30.70 30.60 30.60 30.70
ANLI-R3 31.75 31.17 30.58 30.58 30.33 30.67 30.00 31.00
Average 34.59 35.17 35.02 34.86 35.41 35.20 35.12 35.53
Loss 4.011 3.874 3.966 3.82 3.905 3.776 3.83 3.74
@@
TriviaQA (EM) 1.04 0.79 0.98 1.05 1.16 1.11 1.86 1.21
NaturalQuestions (EM) 0.08 0.17 0.14 0.30 0.30 0.33 0.28 0.39
WebQuestions (EM) 0.59 0.44 0.44 0.64 1.28 0.79 1.33 0.94
LAMBADA 0.16 1.13 0.43 1.28 1.51 1.20 0.49 1.28
HellaSwag 27.77 27.99 27.45 28.29 27.58 28.53 28.86 28.95
StoryCloze 56.33 57.03 57.03 57.40 57.30 58.31 58.63 58.90
WSC 55.44 56.14 56.49 57.89 58.25 58.25 57.54 57.89
WinoGrande 52.01 52.09 50.28 54.30 51.22 52.25 51.54 51.54
Winograd 54.21 56.78 56.78 56.78 61.54 57.14 60.44 61.17
RACE-H 27.93 28.82 27.50 28.47 28.70 29.33 28.73 28.99
RACE-M 33.29 34.05 34.19 33.98 34.54 34.89 33.29 35.58
PIQA 57.13 59.30 56.91 59.85 57.94 59.52 59.52 60.50
ARC-C 22.53 22.10 23.63 22.87 24.06 23.55 24.66 22.95
ARC-E 40.24 40.19 40.19 41.46 41.71 41.71 41.62 42.72
OpenBookQA 30.60 34.20 30.80 33.60 31.00 33.60 34.00 35.00
BoolQ 54.13 52.48 52.45 53.00 55.63 52.57 55.90 52.94
COPA 62.00 65.00 61.00 63.00 61.00 66.00 64.00 64.00
RTE 52.71 50.18 52.35 50.90 50.54 47.65 52.71 48.38
WiC 47.34 47.34 47.34 49.84 47.96 47.96 47.65 47.96
MultiRC 51.44 52.91 52.52 52.62 50.23 53.09 52.41 52.41
RECORD 48.58 50.56 48.99 52.11 50.56 53.21 52.82 54.09
CB 42.86 42.86 42.86 42.86 39.29 44.64 42.86 42.86
ANLI-RI 31.80 32.20 31.80 31.50 32.40 32.50 32.20 32.60
ANLI-R2 30.50 29.40 31.10 29.70 32.00 29.30 30.50 30.30
ANLI-R3 30.08 31.33 30.50 30.67 33.50 30.75 30.67 30.42
Average 36.43 37.02 36.56 37.37 37.24 37.57 37.78 37.75
Loss 3.548 3.484 3.513 3.43 3.456 3.387 3.354 3.348
@@
TriviaQA (EM) 2.09 3.77 2.20 3.87 2.84 4.92 5.18 5.40
NaturalQuestions (EM) 0.11 0.69 0.28 0.72 0.58 0.72 0.91 1.00
WebQuestions (EM) 2.12 2.41 1.08 2.41 1.67 2.71 2.41 3.35
LAMBADA 0.29 1.28 0.66 2.10 1.90 2.66 2.76 3.10
HellaSwag 29.89 30.49 30.05 31.28 31.18 31.87 32.52 32.65
StoryCloze 59.17 60.24 59.54 60.88 60.24 61.79 61.68 62.48
WSC 61.05 59.30 59.30 60.00 61.75 63.16 58.95 62.46
WinoGrande 51.46 52.25 49.57 50.75 52.41 52.72 50.91 52.33
Winograd 55.68 62.27 57.88 64.10 63.00 67.77 61.17 67.40
RACE-H 29.45 29.50 28.90 29.42 29.22 29.87 29.67 29.42
RACE-M 35.31 37.19 36.14 37.26 36.42 37.53 37.60 38.23
PIQA 58.98 60.83 59.58 61.92 59.79 63.00 62.19 62.62
ARC-C 23.38 25.17 23.21 24.06 23.81 24.49 25.00 23.98
ARC-E 42.30 43.86 44.11 45.50 44.53 47.05 46.80 48.36
OpenBookQA 32.80 35.60 34.60 35.80 35.20 37.80 36.80 37.00
BoolQ 53.43 54.56 55.32 53.79 52.87 51.87 54.22 51.31
COPA 61.00 62.00 61.00 62.00 64.00 67.00 60.00 65.00
RTE 52.71 49.10 53.43 49.46 51.62 48.74 54.15 51.26
WiC 47.18 48.75 47.65 47.65 47.65 47.34 47.34 47.34
MultiRC 51.67 51.57 52.70 51.16 53.84 52.58 53.28 52.76
RECORD 54.34 55.93 55.18 56.80 56.75 58.21 58.39 58.97
CB 42.86 44.64 42.86 51.79 42.86 50.00 50.00 44.64
ANLI-RI 32.00 33.20 32.00 33.40 32.50 33.30 32.20 32.40
ANLI-R2 32.60 30.60 30.90 30.10 30.60 30.70 29.80 30.60
ANLI-R3 32.08 31.58 30.75 31.83 32.17 32.00 31.50 31.17
Average 37.75 38.67 37.95 39.12 38.77 40.00 39.41 39.81
Loss 3.316 3.29 3.299 3.235 3.225 3.19 3.16 3.15